OBSERVATION OF Ω = 1/2 STATES IN NiH THROUGH COLLISIONALLY INDUCED FLUORESCENCE

\underline{C. RICHARD^{a}}, P. CROZET, A. J. ROSS, Université Lyon 1; CNRS; LASIM UMR 5579, 43 Bd du 11 novembre 1918, F-69622 Villeurbanne, France; D. W. TOKARYK, Department of Physics and Center for Laser, Atomic, and Molecular Sciences, University of New Brunswick, Fredericton, Canada E3B 5A3.

Fourier transform spectra of collisionally induced fluorescence following isotopically selective laser excitation of NiH at ∼ 550 nm has allowed us to locate an exited Ω′ = 1/2 state of NiH lying 17900 cm^{−1} above the electronic ground state. We identify this as v = 0 of a ^{2}Π_{1/2} state, originating from the Ni^{+} 3d^{8}4s^{1} ^{2}F configuration, from the ab initio studies published by Zou and Liu in 2007^{b}. Emission from the [17.9]0.5 state occurs to v′′ = 0 and 1 of the ^{2}Σ^{+} and ^{2}Π_{1/2} low-lying ligand field states, locating hitherto elusive ^{2}Π_{1/2} f parity levels to within 0.01 cm^{−1}. Collisionally induced fluorescence following laser excitation at lower energies has been recorded in the presence of a magnetic field (0.7 − 1 Tesla), at Doppler resolution. The partially resolved Zeeman patterns have been used to derive effective Landé factors g_{J} for the v = 0 and 1 levels of the low-lying Ω′′ = 5/2 and 3/2 states (^{2}Δ and ^{2}Π states from Ni^{+} 3d^{9}). These are compared with predictions from a revised fit of zero-field energy levels, now including the new observations concerning the Ω′′ = 1/2 states.

aCurrent address: Harvard-Smithsonian Center for Astrophysics, Atomic and Molecular Physics Division, Cambridge MA 02138, USA

bW.L. Zou, W.J. Liu. J. Comp. Chem. 28 (2007) 2286-98